# Effects of Mosaic Crystal Instrument Functions on X-ray Thomson Scattering Diagnostics

### v1

@@ II.3 Intrinsic Rocking Curve
Figure 4: A comparison of the HAPG rocking curve using a single crystal thickness (orange dashed), an uniformly-averaged curve for different thickness (green dashed-dot), and experimental measurements of the rocking curve from Ref. Gerlach et al. (2015) (red points). The experimental curve has a FWHM of 0.00670 (116 $\mu$rad). The theoretical and experimental curves are for a 8.048 keV photon. Note that the theoretical curves have been shifted down in angle by 0.00185 (32.3 $\mu$rad) to center on zero arising from dynamical diffraction theory Zachariasen (1967). Also shown are fits to the experimental data using a Voigt profile (blue solid) and Student’s $t$-distribution (grey dotted). Inset: Focused plot on the wings of the curves. The scale of both axes is shared with the main plot.
Figure 5: A Voigt profile intrinsic rocking curve in terms of the angle of incident on the crystal (left) and the energy on camera (right). The vertical dotted lines indicate cut-offs due to the edges of the crystal. The FWHM of the Gaussian component is 50.4 $\mu$rad (0.00289), while for the Lorentzian component it is 9.48 $\mu$rad (0.000543).
So far, the broadening effects considered have only shown broadening to higher photon energies, yet there is also evidently broadening towards lower photon energies Preston et al. (2020). This broadening comes from the intrinsic rocking curve of the crystallites Zachariasen (1967). The IRC, $W_{\rm RC}(\alpha_{B}-\alpha_{\rm in},\Gamma)$, allows for diffraction to occur even when the Bragg condition is not satisfied, and it is a property of the crystallites themselves. In general, it is a very narrow function. Indeed, for HAPG it is approximately an order of magnitude narrower than the mosaic distribution function Gerlach et al. (2015). Nevertheless, it provides a very important contribution to the crystal instrument function in that it is the only source of broadening towards lower energies. Mechanically, this is because the mosaic broadening in isolation still requires the Bragg condition to be satisfied for a given crystallite, while the IRC allows this condition to be broken. For the IRC in isolation, the reflection rule is still followed, i.e. $\alpha_{\rm out}=\alpha_{\rm in}$, hence there is broadening in both directions (cf. Fig. 5).
@@
The experimental measurement of the IRC, as well as the averaged diffraction pattern, imply that the IRC has a functional form with tails between that of a Lorentzian and a Gaussian – the former being unable to capture the shape near the peak of the IRC, while the latter decays in the tails far too quickly. We note that a recent ray tracing treatment of mosaic crystals considered a rectangular function for the IRC Šmíd, Pan, and Falk (2021); however, we found this function could not reproduce extended tails seen in the scattering data as, like the Gaussian, it decays far too quickly in the tails. Natural choices then for fitting the entire distribution are those which lie between a Gaussian and Lorentzian, such as a Voigt profile (the convolution of a Gaussian and Lorentzian) and the Student’s $t$-distribution. The $t$-distribution is essentially a generalization of the Lorentzian profile to higher powers and connects between the Gaussian and Lorentzian distributions in parameter limits. However, for $|\alpha-\alpha_{B}|>0.008^{\circ}$, the fitted $t$-distribution in Fig. 4 begins to decay faster than the experimental and theoretical curves. Indeed, it was found that the $t$-distribution tended to underpredict the far wings of the broadening towards lower energies in quasi-elastic scattering data. The Voigt profile, also shown in Fig. 4, decays very slowly in the wings. In comparison to the theoretical curves, the Voigt profile captures the slow decay of the wings better than the $t$-distribution. Indeed, the Voigt profile was found to be able to reliably fit the wings of the quasi-elastic peak even to very low energies. We therefore choose to approximate the IRC using the Voigt profile (scipy.special.voigt_profile from the SciPy package for Python Virtanen et al. (2020)), where the Lorentzian and Gaussian FWHMs are left as free parameters in the fitting. While this choice is motivated by the need for a curve that provides good fits to the quasi-elastic peak, evidently it would be preferable to have additional measurements of the HAPG IRC to higher rocking angles so that the decay of the wings could be better modelled, either with an approximate function as done here or with diffraction theory.
@@
Like the mosaicity, the IRC also depends on the photon energy, with photons of higher energy observing an IRC with a smaller width Gerlach et al. (2015); Zachariasen (1967), though like the mosaicity the IRC width depends quite weakly on photon energy. An additional complication is that the shape of the IRC of a crystallite also depends on the photon energy Zachariasen (1967). But, for the ranges on the detectors of interest here, both of these dependencies are assumed to be sufficiently weak that the IRC can be approximated as a Voigt profile across the entire range of the detector.
@@
In Fig. 5, an example Voigt profile IRC is plotted both in terms of the angular difference and the energy on camera, but assuming the width is fixed. In terms of the angle of incidence, the curves have the same shape except for the position of the cut-offs due to the crystal edges. This is complicated when the mosaicity is introduced as the IRC now allows for the Bragg condition to be broken for the crystallites. As a result, for a given angle of incidence there are now a number of crystallites which a photon can reflect off. For the present work, at each angle of incidence on the crystal, we multiply the mosaic distribution function by the IRC to account for the probability of encountering a crystallite at a given angle from the surface normal, and the reduction in reflectivity from being away from the Bragg condition. In terms of energy observed on camera, we observe some differences in the IRC due to the different positions of the central Bragg angle [cf. Eq. (10)]. In other words, even keeping the angular IRC the same for a fixed photon energy, the geometry of the setup still results in variations in the broadening.
@@ II.4 Model Application
Figure 6: A comparison of the crystal instrument for different maximum reflection orders in 40 $\mu$m (left) and 100 $\mu$m (right) of HAPG. The green dashed-dot curve shows the IF when neglecting depth broadening. For the mosaicities considered here, for the 40 $\mu$m crystal shows little difference in the broadening from only $p=1$ reflections (red solid) and when including up to $p=4$ (blue dashed) reflections, while for the 100 $\mu$m crystal the multiple reflections have some noticeable contribution to the wings of the IF.
In this work, we examine the instrument broadening of mosaic HAPG crystals with parameters similar to those available at the European XFEL (EuXFEL) in Germany Preston et al. (2020); Zastrau et al. (2021). Specifically, we consider the crystals that have a radius of curvature of 80 mm, a length of 80 mm, and thicknesses of 40 $\mu$m and 100 $\mu$m. In the present mosaic model, the source is treated as a point at the origin; the crystal is a line at a fixed distance below the source $r$, but can be varied in size and distance along the dispersion axis; and the camera is likewise a line parallel to the crystal that can be varied in height and distance from the source; see Fig. 1. As mentioned previously, this explicitly neglects any broadening effects off the dispersion axis on the camera and along the bent crystal. There are also additional effects, e.g. the strain within the crystallites or from the bending process Legall et al. (2006a), that are not explicitly accounted for in the model. Nevertheless, this simplified treatment of the spectrometer still provides good agreement with experimental measurements.
In order to converge the shapes of the curves, 450 angles of incident on the crystal were sampled, and the IRC was linearly sampled across 3750 angles between ±$\pm 1.5^{\circ}$. For the $G_{p}$ term in Eq. (7), we considered up to $p=4$ and 30 reflection layers. While it is obvious that depth effects must be considered, multiple reflections within the crystal did not change the IF significantly, as can be observed in the Fig. 6. For the 40 $\mu$m crystal, there was no observable difference between the IF whether considering a maximum order of $p=1$ or $p=4$. For the 100 $\mu$m crystal, slight differences in the IF can be observed in the far wings.
@@
The procedure to calculate the crystal reflectivity is as follows: the extent and position of the crystal and camera compared to the source are defined, and the photon energy $E_{0}$ is also chosen. Additionally, pixels on the camera are defined for later binning the instrument function. The crystal is then divided into a large number of sections with co-ordinates $(x_{\rm crys},r)$ which are then associated to different angles of incidence $\alpha_{\rm in}$. For each $\alpha_{\rm in}$, the differential reflectivity versus $z_{\rm eff}$ is calculated using Eq. (7). To determine the crystal broadening observed on camera, the shift within the crystal is first estimated to be $\Delta x_{\rm crys}=z_{\rm eff}/\tan(\alpha_{\rm in})$, as is done in Ref. Schlesiger et al. (2017) – that is, a photon is treated as travelling its full effective depth into the crystal before it encounters the crystallite it reflects off. It is also at this depth that the effect of the IRC is applied. The final position of the photon on the camera axis is then traced and binned into the camera pixels to develop the instrument function as observed by the camera, which allows for direct comparison with experimental measurements of the instrument function.
@@ III.1 SIF Model versus Experiment
Figure 7: A comparison of the crystal IF model (red dashed) fitted to experimental measurements (black) of the quasi-elastic signal of a monochromated beam with (a) $E_{0}=7.7$ keV on 50 $\mu$m thick Al with 40 $\mu$m HAPG, (b) $E_{0}=8.5$ keV on 8 $\mu$m thick polypropylene (PP) with 40 $\mu$m HAPG, and (c) $E_{0}=8.5$ keV on 8 $\mu$m thick PP with 100 $\mu$m HAPG. The angles in the brackets are the scattering angles at which the spectra were taken. The grey shaded regions indicate the uncertainty in the experimental spectra. The theoretical SIFs are plotted on energy axes that extended beyond the experimental spectral range of the camera.
@@
We start by comparing the HAPG crystal IF fit using the model on experimental measurements of the IF in order to verify the applicability of the model. To do this, we compare the IF predicted by the model to the quasi-elastic scattering signal from XRTS measurements. Except for very high resolution setups McBride et al. (2018); Wollenweber et al. (2021); Descamps et al. (2022), the quasi-elastic feature is essentially a delta function and so it reproduces the SIF, assuming Eq. 5 is a convolution. A material that produces no inelastic scattering would be ideal as it would give direct access to the SIF. In practice, we must settle with using materials that have very weak inelastic scattering in the vicinity of the quasi-elastic feature for the particular scattering vector examined. Plastics such as PMMA and polypropylene (PP) have very prominent quasi-elastic features relative to the inelastic scattering, while Al in backscattering geometry has inelastic features that are very far from the quasi-elastic signal. In combination with a monochromated beam that allows the source contribution to the SIF to be neglected, here we will use PP and Al to measure the spectrometer instrument function for experiments that were performed at the EuXFEL.
@@
The experimental data presented in this section was collected as part of two different experiments at the EuXFEL. While the primary goals of the experiments were not to measure the instrument functions of the HAPG crystals, the experiments nevertheless provided high quality data of the quasi-elastic feature. In both cases, the FEL beam was self-seeded and then passed through a four-bounce Si (111) monochromator with an acceptance range of 0.8 eV. The monochromator has been shown to result in beams with FWHM $\sim 500$ meV Gawne et al. (2024). The broadening of the signal by the beam is therefore negligible, which was verified by broadening the modelled IF by a 500 meV and not observing any difference in the spectrum. We may therefore focus on the broadening by the crystal itself. In all cases, the XRTS signal was collected in ambient conditions – there is therefore no upshfited part to the DSF allowing for detailed examination of the broadening up to high energies above the quasi-elastic signal. Finally, the samples were thin, so that broadening due to scattering from different points across the sample depth were found to be negligible.
@@
The first measurement we compare against is on 50 $\mu$m Al at a scattering angle of $167^{\circ}$ and a photon energy of 7.7 keV, plotted in Fig. 7 (a). The spectrum was collected using a 40 $\mu$m HAPG crystal. The first inelastic feature is $\sim 72$ eV below the quasi-elastic feature, allowing for the broadening towards lower energies to be examined without much influence from the inelastic scattering. We find good agreement between the model and the experimental data. Notably, the slow decay of the quasi-elastic peak to lower energies further supports the notion that the IRC decays in a manner between that of a Gaussian and Lorentzian. As for the upshifted broadening, we also find good agreement in the spectral region that was measured, however the spectrum cuts off at 34 eV above the elastic, and this limited range in the upshifted part of the spectrum makes it difficult to assess whether this tail has been accurately modelled. Still, both the experimental data and the model suggest that the quasi-elastic signal is substantial and comparable to the detectable signal level – $2.5\times 10^{-4}$ weaker than the elastic peak in this dataset – even $>100$ eV away from the elastic peak. Therefore, in order to meaningfully measure the full instrument function of the spectrometer, the elastic peak needs to be positioned relatively far from the camera edges.
@@
The second measurement of the crystal IF is on 8 $\mu$m PP with a photon energy of 8.5 keV in both backwards (166) and forwards (17) scattering, plotted in Fig. 7 (b) and (c), respectively. The backscattering data was collected on the 40 $\mu$m thick crystal while the forward scattering data was collected on the 100 $\mu$m thick crystal. Like Al, PP has relatively weak inelastic scattering, though now it is much closer to the quasi-elastic signal than in Al. As a result, the downshifted broadening is harder to distinguish from the elastic peak. However the main shape of the quasi-elastic peak and the upshifted wing is unaffected by the broadened inelastic scattering contributions as the inelastic scattering is orders of magnitude weaker than the elastic feature. In this dataset the elastic scattering was position 150–200 eV below the edge of the detector so that the upshifted broadening can be examined. The experimental data indicates that there is substantial signal in the wings even some 200 eV away from the quasi-elastic peak. In both spectra, we find the model is able to reproduce the main peak of the quasi-elastic feature. For the 40 $\mu$m crystal, the model reproduces the upshifted broadening quite well. Above 8.53 keV, the spectrum becomes quite noisy, making it difficult to assess the quality of the model. Above 8.6 keV, the noise level of the data is reached, which may explain the gaps in the spectrum. More shots would be required to unambiguously assess the quality of the model.
Fig. 7 Label $T_{c}$ ($\mu$m) $E_{0}$ (keV) $\gamma$ (deg.) $F_{L}$ (deg.) $F_{G}$ (deg.) $F_{V}$ (deg.)
(a) 40 7.7 0.127 $1.40\times 10^{-3}$ $5.20\times 10^{-3}$ $5.99\times 10^{-3}$
(b) 40 8.5 0.075 $0.54\times 10^{-3}$ $2.89\times 10^{-3}$ $3.19\times 10^{-3}$
(c) 100 8.5 0.078 $0.56\times 10^{-3}$ $3.50\times 10^{-3}$ $3.81\times 10^{-3}$
@@
The final fit parameters for the mosaicity $\gamma$, and FWHM for the Lorentzian $F_{L}$ and Gaussian $F_{G}$ components of the IRC are given in Table 1. Note that due to the depth broadening being akin to an exponential, it results in a small shift of the maximum away from the elastic peak by $\sim 1.6$ eV, which is corrected for before the modelled and experimental spectra are compared to evaluate the fit. The three fitted mosaicities are all $\sim 0.1^{\circ}$, which is reasonable for HAPG Gerlach et al. (2015). In all three cases, the Gaussian component of the IRC is larger than the Lorentzian component, but the presence of the latter means the tails will be extended to high angles and at large energies shifts.
Also provided in Table 1 are the FWHM of the Voigt profiles $F_{V}$ using Kielkopf’s approximation Kielkopf (1973). The two widths from the PP data are quite similar, but narrower than the measurements of the IRC on Al. Measurements of the IRC presented in Ref. Gerlach et al. (2015) suggest that, at the two photon energies examined here, the IRC of HAPG would be $\sim 0.0060\pm 0.0003^{\circ}$. The PP measurements are substantially narrower than this, and fixing the IRC width to $0.0060^{\circ}$ leads to an overestimation of the main peak of the quasi-elastic feature.
In all three cases, we find that the the instrument functions are highly asymmetric, as is shown in Fig. 8. Near to the central photon energy, the broadening is strongly asymmetric towards the upshifted side of the spectrum. Once the mosaicity drops off as the edges of the crystal are reached at very high energy shifts, the broadening is then strongly asymmetric towards the downshifted side of the spectrum. However, the intensity is so low at such large energy shifts it may not be detectable. Additionally, such large energy shifts may also fall outside the spectral range of the camera. In general then, the bulk of the broadening will be towards higher energies. In the context of XRTS, where the ratio of the upshifted to downshifted intensity is related via the detailed balance relation, the asymmetry of the SIF will lead to an enhancement of the upshifted side of the spectrum. If a symmetric SIF is then used to infer temperature in forward modelling, or in the deconvolution step of in the imaginary-time domain, then detailed balance a higher temperature may be inferred in order to compensate for the asymmetry of the actual SIF.

### v2

@@ II.3 Intrinsic Rocking Curve
Figure 4: A comparison of the HAPG rocking curve using a single crystal thickness (orange dashed), an uniformly-averaged curve for different thickness (green dashed-dot), and experimental measurements of the rocking curve from Ref. Gerlach et al. (2015) (red points). The experimental curve has a FWHM of 116 $\mu$rad. The theoretical and experimental curves are for a 8.048 keV photon. Note that the theoretical curves have been shifted down in angle by 32.3 $\mu$rad to center on zero arising from dynamical diffraction theory Zachariasen (1967). Also shown are fits to the experimental data using a Voigt profile (blue solid) and Student’s $t$-distribution (grey dotted). Inset: Focused plot on the wings of the curves. The scale of both axes is shared with the main plot.
Figure 5: A Voigt profile intrinsic rocking curve in terms of the angle of incidence on the crystal (left) and the energy on the detector (right). The vertical dotted lines indicate cut-offs due to the edges of the crystal. The FWHM of the IRC is 55.7 $\mu$rad, with a Gaussian FWHM of 50.4 $\mu$rad and a Lorentzian FWHM of 9.48 $\mu$rad.
So far, the broadening effects considered have only shown broadening to higher photon energies, yet there is also evidently broadening towards lower photon energies Preston et al. (2020). This broadening comes from the intrinsic rocking curve (IRC) of the crystallites Zachariasen (1967). The IRC, $W_{\rm RC}(\alpha_{B}-\alpha_{\rm in},\Gamma)$, allows for diffraction to occur even when the Bragg condition is not satisfied, and it is a property of the crystallites themselves. In general, it is a very narrow function. Indeed, for HAPG it is approximately an order of magnitude narrower than the mosaic distribution function Gerlach et al. (2015). Nevertheless, it provides a very important contribution to the crystal instrument function in that it is the only source of broadening towards lower energies. Mechanically, this is because the mosaic broadening in isolation still requires the Bragg condition to be satisfied for a given crystallite, while the IRC allows this condition to be broken. For the IRC in isolation, the reflection rule is still followed, i.e. $\alpha_{\rm out}=\alpha_{\rm in}$, hence there is broadening in both directions (cf. Fig. 5).
@@
The experimental measurement of the IRC, as well as the averaged theoretical curve, imply that the IRC has a functional form with tails between that of a Lorentzian and a Gaussian – the former being unable to capture the shape near the peak of the IRC, while the latter decays in the tails far too quickly. We note that a recent ray tracing treatment of mosaic crystals considered a rectangular function for the IRC Šmíd, Pan, and Falk (2021); however, we found this function could not reproduce extended tails seen in the scattering data as, like the Gaussian, it decays far too quickly in the tails. Natural choices then for fitting the entire distribution are those which lie between a Gaussian and Lorentzian, such as a Voigt profile (the convolution of a Gaussian and Lorentzian) and the Student’s $t$-distribution. The $t$-distribution is essentially a generalization of the Lorentzian profile to higher powers and connects between the Gaussian and Lorentzian distributions in parameter limits. However, for $|\alpha-\alpha_{B}|>0.008^{\circ}$, the fitted $t$-distribution in Fig. 4 begins to decay faster than the experimental and theoretical curves. Indeed, it was found that the $t$-distribution tended to underpredict the far wings of the broadening towards lower energies in quasi-elastic scattering data. The Voigt profile, also shown in Fig. 4, decays very slowly in the wings. In comparison to the theoretical curves, the Voigt profile captures the slow decay of the wings better than the $t$-distribution. Indeed, the Voigt profile was found to be able to reliably fit the wings of the quasi-elastic peak, even to large energy shifts below the peak. We therefore choose to approximate the IRC using the Voigt profile (scipy.special.voigt_profile from the SciPy package for Python Virtanen et al. (2020)), where the Lorentzian and Gaussian FWHMs are left as free parameters in the fitting. While this choice is motivated by the need for a curve that provides good fits to the quasi-elastic peak, evidently it would be preferable to have additional measurements of the HAPG IRC to higher rocking angles so that the decay of the wings could be better modelled, either with an approximate function as done here or with diffraction theory.
@@
Like the mosaicity, the IRC also depends on the photon energy, with photons of higher energy observing an IRC with a smaller width Gerlach et al. (2015); Zachariasen (1967), though like the mosaicity the IRC width depends quite weakly on photon energy. An additional complication is that the shape of the IRC of a crystallite also depends on the photon energy Zachariasen (1967). But, for the photon energies of interest here (6.90-7.7 keV, 7.8-8.7 keV), both of these dependencies are assumed to be sufficiently weak that the IRC can be approximated as a Voigt profile across the entire detection range.
@@
In Fig. 5, an example Voigt profile IRC is plotted both in terms of the angular difference and the energy on the detector, but assuming the width is fixed. In terms of the angle of incidence, the curves have the same shape except for the position of the cut-offs due to the crystal edges. This is complicated when the mosaicity is introduced as the IRC now allows for the Bragg condition to be broken for the crystallites. As a result, for a given angle of incidence there are now a number of crystallites which a photon can reflect off. For the present work, at each angle of incidence on the crystal, we multiply the mosaic distribution function by the IRC to account for the probability of encountering a crystallite at a given angle from the surface normal, and the reduction in reflectivity from being away from the Bragg condition. In terms of energy observed on the detector, we observe some differences in the IRC due to the different positions of the central Bragg angle [cf. Eq. (10)]. In other words, even keeping the angular IRC the same for a fixed photon energy, the geometry of the setup still results in variations in the broadening.
@@ II.4 Model Application
Figure 6: A comparison of the crystal instrument functions for different maximum reflection orders in 40 $\mu$m (left) and 100 $\mu$m (right) of HAPG. The green dashed-dot curve shows the IF when neglecting depth broadening. For the mosaicities considered here, for the 40 $\mu$m crystal shows little difference in the broadening from only $p=1$ reflections (red solid) and when including up to $p=4$ (blue dashed) reflections, while for the 100 $\mu$m crystal the multiple reflections have some noticeable contribution to the wings of the IF.
In this work, we examine the instrument broadening of mosaic HAPG crystals with parameters similar to those available at the European XFEL (EuXFEL) in Germany Preston et al. (2020); Zastrau et al. (2021). Specifically, we consider the crystals that have a radius of curvature of 80 mm, a length of 80 mm, and thicknesses of 40 $\mu$m and 100 $\mu$m. In the present mosaic model, the source is treated as a point at the origin; the crystal is a line at a fixed distance below the source $r$, but can be varied in size and distance along the dispersion axis; and the detector is likewise a line parallel to the crystal that can be varied in height and distance from the source; see Fig. 1. As mentioned previously, this explicitly neglects any broadening effects off the dispersion axis on the detector by taking a narrow region of interest around the focus line, and along the bent crystal because the crystal is narrow at only 30 mm. There are also additional effects, e.g. the strain within the crystallites or from the bending process Legall et al. (2006a), that are not explicitly accounted for in the model. Nevertheless, this simplified treatment of the spectrometer still provides good agreement with experimental measurements.
In order to converge the shapes of the curves, 450 angles of incidence on the crystal were sampled, and the IRC was linearly sampled across 3750 angles between ±$\pm 1.5^{\circ}$. For the $G_{p}$ term in Eq. (7), we considered up to $p=4$ and 30 reflection layers. While it is obvious that depth effects must be considered, multiple reflections within the crystal did not change the IF significantly, as can be observed in the Fig. 6. For the 40 $\mu$m crystal, there was no observable difference between the IF whether considering a maximum order of $p=1$ or $p=4$. For the 100 $\mu$m crystal, slight differences in the IF can be observed in the far wings.
@@
The procedure to calculate the crystal reflectivity is as follows: the extent and position of the crystal and detector compared to the source are defined, and the energy of the incident photon $E_{0}$ is also chosen. Additionally, pixels on the detector are defined for later binning the instrument function. The crystal is then divided into a large number of sections with co-ordinates $(x_{\rm crys},r)$ which are then associated to different angles of incidence $\alpha_{\rm in}$. For each $\alpha_{\rm in}$, the differential reflectivity versus $z_{\rm eff}$ is calculated using Eq. (7). To determine the crystal broadening observed on the detector, the shift within the crystal is first estimated to be $\Delta x_{\rm crys}=z_{\rm eff}/\tan(\alpha_{\rm in})$, as is done in Ref. Schlesiger et al. (2017) – that is, a photon is treated as travelling its full effective depth into the crystal before it encounters the crystallite it reflects off. It is also at this depth that the effect of the IRC is applied. The final position of the photon on the detector axis is then traced and binned into the detector pixels to develop the instrument function as observed by the detector, which allows for direct comparison with experimental measurements of the instrument function.
@@ III.1 SIF Model versus Experiment
Figure 7: A comparison of the crystal IF model (red dashed) fitted to experimental measurements (black) of the quasi-elastic signal of a monochromated beam with (a) $E_{0}=7.7$ keV on 50 $\mu$m thick Al with 40 $\mu$m HAPG, (b) $E_{0}=8.5$ keV on 8 $\mu$m thick polypropylene (PP) with 40 $\mu$m HAPG, and (c) $E_{0}=8.5$ keV on 8 $\mu$m thick PP with 100 $\mu$m HAPG. The angles in the brackets are the scattering angles at which the spectra were taken. The grey shaded regions indicate the uncertainty in the experimental spectra. The theoretical SIFs are plotted on energy axes that extended beyond the experimental spectral range of the detector.
@@
We start by comparing the HAPG crystal IF fit using the model on experimental measurements of the IF in order to verify the applicability of the model. To do this, we compare the IF predicted by the model to the quasi-elastic scattering signal from XRTS measurements. Except for very high resolution setups McBride et al. (2018); Wollenweber et al. (2021); Descamps et al. (2022), the quasi-elastic feature is essentially a delta function and so it reproduces the SIF, assuming Eq. 5 is a convolution. A material that produces no inelastic scattering would be ideal as it would give direct access to the SIF. In practice, we must settle with using materials that have very weak inelastic scattering in the vicinity of the quasi-elastic feature for the particular scattering vector examined. Note that this means there is inevitably some uncertainty in the IF in the vicinity of the inelastic features. Plastics such as PMMA and polypropylene (PP) have very prominent quasi-elastic features relative to the inelastic scattering, while Al in backscattering geometry has inelastic features that are very far from the quasi-elastic signal. In combination with a monochromated beam that allows the source contribution to the SIF to be neglected, here we will use PP and Al to measure the spectrometer instrument function for experiments that were performed at the EuXFEL.
@@
The experimental data presented in this section was collected as part of two different experiments at the EuXFEL. While the primary goals of the experiments were not to measure the instrument functions of the HAPG crystals, the experiments nevertheless provided high quality data of the quasi-elastic feature. In both cases, the FEL beam was self-seeded and then passed through a four-bounce Si (111) monochromator with an acceptance range of 0.8 eV. The monochromator has been shown to result in beams with FWHM $\sim 500$ meV Gawne et al. (2024). The broadening of the signal by the beam is therefore negligible, which was verified by broadening the modelled IF by a 500 meV and not observing any difference in the spectrum. Additionally, in both experiments the spot sizes on the targets were 10 $\mu$m. In von Hámos geometry, there is a one-to-one imaging of the spot on the detector, so the finite spot size broadening in the dispersive direction is smaller than the pixel size of the detectors used. The broadening by these small spot sizes was therefore also found to be negligible. We may therefore focus on the broadening by the crystal itself. In all cases, the XRTS signal was collected in ambient conditions – there is therefore no upshfited part to the DSF allowing for detailed examination of the broadening up to high energies above the quasi-elastic signal. The thickness of the HAPG crystals of interest have a standard deviation of 2 $\mu$m, so the broadening due to crystal surface roughness is negligible and not included in the model. Finally, the samples were thin, so that broadening due to scattering from different points across the sample depth were found to be negligible.
@@
The first measurement we compare against is on 50 $\mu$m Al at a scattering angle of $167^{\circ}$ and a photon energy of 7.7 keV, plotted in Fig. 7 (a). The spectrum was collected using a 40 $\mu$m HAPG crystal. The first inelastic feature is $\sim 72$ eV below the quasi-elastic feature, allowing for the broadening towards lower energies to be examined without much influence from the inelastic scattering. We find good agreement between the model and the experimental data. Notably, the slow decay of the quasi-elastic peak to lower energies further supports the notion that the IRC decays in a manner described by a Voigt function. As for the upshifted broadening, we also find good agreement in the spectral region that was measured, however the spectrum is only detected up to 34 eV above the elastic, and this limited range in the upshifted part of the spectrum makes it difficult to assess whether this tail has been accurately modelled. Still, both the experimental data and the model suggest that the quasi-elastic signal is substantial and comparable to the detectable signal level – $2.5\times 10^{-4}$ weaker than the elastic peak in this dataset – even $>100$ eV away from the elastic peak. Therefore, in order to meaningfully measure the full instrument function of the spectrometer, the elastic peak needs to be positioned relatively far from the detector edges.
@@
The second measurement of the crystal IF is on 8 $\mu$m PP with a photon energy of 8.5 keV in both backwards (166) and forwards (17) scattering, plotted in Fig. 7 (b) and (c), respectively. The backscattering data was collected on the 40 $\mu$m thick crystal while the forward scattering data was collected on the 100 $\mu$m thick crystal. Like Al, PP has relatively weak inelastic scattering, although it is much closer to the quasi-elastic signal than in Al. As a result, the downshifted broadening is harder to distinguish from the elastic peak. However the main shape of the quasi-elastic peak and the upshifted wing is unaffected by the broadened inelastic scattering contributions as the inelastic scattering is orders of magnitude weaker than the elastic feature. In this dataset the elastic scattering was positioned 150–200 eV below the high-energy edge of the detector so that the upshifted broadening can be examined. The experimental data indicates that there is substantial signal in the wings even some 200 eV away from the quasi-elastic peak. In both spectra, we find the model is able to reproduce the main peak of the quasi-elastic feature. For the 40 $\mu$m crystal, the model reproduces the upshifted broadening quite well. Above 8.53 keV, the spectrum becomes quite noisy, making it difficult to assess the quality of the model. Above 8.6 keV, the noise level of the data is reached, which explains the gaps in the spectrum as the data is collected in the single photon regime. More shots would be required to unambiguously assess the quality of the model in this region.
Fig. 7 Label $T_{c}$ ($\mu$m) $E_{0}$ (keV) $\gamma$ (deg.) $F_{L}$ ($\mu$rad) $F_{G}$ ($\mu$rad) $F_{V}$ ($\mu$rad)
(a) 40 7.7 0.067 $23.4$ $83.8$ $97.0$
(b) 40 8.5 0.075 $9.42$ $50.4$ $55.7$
(c) 100 8.5 0.078 $9.77$ $61.1$ $66.5$
@@
The final fit parameters for the mosaicity $\gamma$, and FWHM for the Lorentzian $F_{L}$ and Gaussian $F_{G}$ components of the IRC are given in Table 1. Note that due to the depth broadening being akin to an exponential, it results in a small shift of the maximum away from the elastic peak by $\sim 1.6$ eV, which is corrected for before the modelled and experimental spectra are compared to evaluate the fit. The three fitted mosaicities are all $\lesssim 0.1^{\circ}$, which is expected for HAPG Gerlach et al. (2015). In all three cases, the Gaussian component of the IRC is larger than the Lorentzian component, but the presence of the latter means the tails will be extended to high angles and at large energies shifts.
Also provided in Table 1 are the FWHM of the Voigt profiles $F_{V}$ using Kielkopf’s approximation Kielkopf (1973). The two widths from the PP data are quite similar, but narrower than the measurements of the IRC on Al. Measurements of the IRC presented in Ref. Gerlach et al. (2015) suggest that, at the two photon energies examined here, the IRC of HAPG would be $\sim 104\pm 5$ $\mu$rad. The PP measurements are substantially narrower than this, and fixing the IRC width to 104 $\mu$rad leads to an overestimation of the main peak of the quasi-elastic feature.
In all three cases, we find that the the instrument functions are highly asymmetric, as is shown in Fig. 8. Near to the central photon energy, the broadening is strongly asymmetric towards the upshifted side of the spectrum. Once the mosaicity drops off as the edges of the crystal are reached at very high energy shifts, the broadening is then strongly asymmetric towards the downshifted side of the spectrum. However, the intensity is so low at such large energy shifts it may not be detectable. Additionally, such large energy shifts may also fall outside the spectral range of the detector. In general then, the bulk of the broadening will be towards higher energies. In the context of XRTS, where the ratio of the upshifted to downshifted intensity is related via the detailed balance relation, the asymmetry of the SIF will lead to an enhancement of the upshifted side of the spectrum. If a symmetric SIF is then used to infer temperature in forward modelling, or in the deconvolution step of in the imaginary-time domain, then detailed balance a higher temperature may be inferred in order to compensate for the asymmetry of the actual SIF.